# Weakly Supervised Gaussian Contrastive Grounding with Large Multimodal Models for Video Question Answering

### v1

@@ Abstract
Video Question Answering (VideoQA) aims to answer natural language questions based on the information observed in videos. Despite the recent success of Large Multimodal Models (LMMs) in image-language understanding and reasoning, they deal with VideoQA insufficiently by simply taking uniformly sampled frames as visual inputs, which ignores question-relevant visual clues. Moreover, there are no human annotations for question-critical timestamps in existing VideoQA datasets. In light of this, we propose a novel weakly supervised framework to enforce the LMMs to reason out the answers with question-critical moments as visual inputs. Specifically, we fuse the question and answer pairs as event descriptions to find multiple keyframes as target moments, which will be pseudo-labels. With these pseudo-labels as additionally weak supervision, we devise a lightweight Gaussian-based Contrastive Grounding (GCG) module. GCG learns multiple Gaussian functions to characterize the temporal structure of the video, and sample question-critical frames as positive moments to be the visual inputs of LMMs. Extensive experiments on several VideoQA benchmarks verify the effectiveness of our framework, and we achieve substantial improvements compared to previous state-of-the-art methods
✉ Corresponding AuthorCodes will be released after review
@@
Figure 1: We argue that the information in uniformly sampled frames is insufficient for LMMs to answer the question correctly. Therefore, we utilize the fused event description to provide additional weak supervision and generate weight distributions for each video moment. We align the positive description-moment pairs while pushing away negative ones. In this way, we can effectively select question-critical moments for LMMs to reason out the answer.
1 Introduction
@@
Video Question Answering (VideoQA) stands at the forefront of developing intelligent systems that can reason about causal and temporal relations and answer natural language questions in videos, which is an essential manifestation of human intelligence. Despite significant advancements in recent years, VideoQA remains a challenging problem that requires models to comprehensively understand and dynamically align the semantics of both the visual world and natural language.
@@
With the notable progress in vision-language pre-training techniques Radford et al. (2021); Li et al. (2022b) and Large Language Models (LLMs) Touvron et al. (2023); Brown et al. (2020), Large Multimodal Models (LMMs) Li et al. (2023c); Dai et al. (2023); Liu et al. (2023) have showcased impressive capabilities across various image-language tasks. However, limited by the long sequence frames in videos and computation costs, current LMMs fall short when applied to VideoQA. They simply concatenate the visual tokens of uniformly/randomly sampled, sparse frames (e.g., 4 frames) as the visual inputs for answer prediction. Such a uniform sampling strategy does not consider the specific question, treating all frames equally and introducing redundancy, particularly when tackling complicated questions within longer videos.
@@
Therefore, it’s necessary to localize the moments crucial for answering the question for LMMs (as the positive moments shown in Figure 1). Notably, different from the task of Temporal Sentence Grounding (TSG) Gao et al. (2017); Zhang et al. (2023) which aims to localize a video moment described by a declarative sentence, the grounding mechanism in VideoQA features some unique challenges. First, questions in VideoQA are interrogative sentences, and they lack explicit information about the answer content needed to be grounded. For instance, there is a semantic gap between the interrogative question [How does the boy react after opening the present?] and the declarative description [The boy carries up the dog after opening the present.]. Thus, models are required to both localize the moment [after opening the present] explicitly shown in the question and identify the implicit answer moment [carry up the dog], demanding the causal-temporal reasoning. Second, VideoQA aims to correctly answer the questions of videos, rather than solely grounding specific video moments, and there are no human annotations for the timestamps of question-critical moments in existing VideoQA datasets.
@@
To address these challenges, we introduce a weakly supervised framework by discovering question-critical moments with Gaussian-based Contrastive Grounding (GCG). As labeling the timestamps of question-critical moments is labor-intensive and subjective, we leverage the visual-language alignment capability of the CLIP models Fang et al. (2023) to weakly provide timestamps of question-critical moments. In detail, we fuse the textual question and answer to generate a declarative sentence as the event description, and then compute the similarities between the description and each frame. Frames with the highest scores will be the keyframes of target moments. We observe that LMMs with these keyframes as visual inputs showcased significant improvements in VideoQA datasets. With these pseudo-labeled timestamps as additional weak supervision, our GCG will learn multiple Gaussian functions to characterize the inherent temporal structure of the video, and distinguish the positive video moments (green in Figure 1) from negative video moments (orange and yellow in Figure 1). The positive moments are crucial for answering the question and will be the visual inputs of LMMs for final answer prediction. Moreover, to ensure the selected positive moments are closest to the event description, our GCG also includes a contrastive objective He et al. (2020) that can align the positive description-moment pairs while pushing away negative ones. Different from previous works like SeViLA Yu et al. (2023) which pre-train an additional LMM as the keyframe localizer with other TSG datasets like QV-Highlights Lei et al. (2021), our method is lightweight for end-to-end training and flexible for LMMs.
To summarize, we make the following contributions:
  • We propose a weakly supervised grounding framework for VideoQA, by utilizing the alignment capability of the CLIP models to provide pseudo-labeled timestamps of keyframes without human-labor annotated costs.
  • We devise the Gaussian-based Contrastive Grounding (GCG) for weakly-grounded selection of question-critical moments, enhancing the efficiency and interpretability of LMMs when applied to VideoQA.
  • We conduct extensive experiments to verify the effectiveness of our proposed method, and achieve substantial improvements on six challenging VideoQA benchmarks including NExT-QA, Causal-VidQA, Intent-QA, ActivityNet-QA, MSVD-QA, and MSRVTT-QA.
2 Related Works
2.1 Large Multimodal Models (LMMs)
@@
LMMs Dai et al. (2023); Li et al. (2023c); Ye et al. (2023); Liu et al. (2023); Alayrac et al. (2022) in their current form primarily function as image-to-text generative models, taking images as input and generating text sequences. These models have demonstrated strong capabilities in image-language understanding and reasoning by adapting frozen language models to frozen image encoders with trainable connection modules, following large-scale image-text pretraining. The connection module can either be a transformer-based architecture like Q-former in InstructBLIP and BLIP-2 Dai et al. (2023); Li et al. (2023c), Perceiver Resampler in Flamingo Alayrac et al. (2022), or a simple linear layer in LLaVA Liu et al. (2023). Most current LMMs are essentially image-based models, and they simply concatenate the visual tokens extracted from uniformly sampled, sparse frames as visual inputs for video-language tasks. This results in a lack of temporal modeling ability and emphasizes the necessity of selecting specific video moments, particularly for addressing the demands of reasoning-based VideoQA tasks. In this paper, our goal is to enhance the causal-temporal reasoning abilities of LMMs without additional pretraining on video-text corpora, by discovering the question-critical moments with our weakly-supervised Gaussian-based Contrastive Grounding.
2.2 Temporal Grounding in VideoQA
@@
Early VideoQA benchmarks Yu et al. (2019); Xu et al. (2017); Jang et al. (2017) focus on descriptive questions (e.g., [what’s the man doing]) within short video clips, rarely going beyond a recognition of the objects and actions. Instead, more recent VideoQA benchmarks Li et al. (2022a); Xiao et al. (2021); Li et al. (2023b) like NExT-QA emphasize counterfactual, temporal, and causal reasoning involving multiple entities and relations, demanding the ability to uncover the causes behind specific events within longer videos, necessitating localizing a text query to specific moments.
@@
In light of this, ATP Buch et al. (2022) utilizes the tool of atemporal probe to select a single frame without temporal information for downstream tasks. MIST Gao et al. (2023) and TranSTR Li et al. (2023e) fuse frames with the mechanism of adaptive temporal rationalization and iterative spatial-temporal attention, respectively. NExT-GQA Xiao et al. (2023) constructs grounding labels in the test set of NExT-QA and uses a single Gaussian mask to fuse frames along the temporal dimension. SeViLA Yu et al. (2023), similar to us, utilizes the LMM (BLIP-2) for VideoQA. However, SeViLA uses two LMMs to generate pseudo-labels and answer questions respectively, with extra pre-training on TSG datasets Lei et al. (2021) and a multi-stage training scheme. Different from previous works, we utilize the CLIP Radford et al. (2021); Fang et al. (2023) models to automatically provide weak supervision for grounding, and our lightweight GCG module learns multiple Gaussian functions to generate both positive and negative moments in an end-to-end manner, with an additionally contrastive objective to distinguish positive ones from negative ones for frame selection.
3 Preliminary: LMMs for VideoQA
@@
We take InstructBLIP Dai et al. (2023) as an example to illustrate how LMMs deal with VideoQA. LMMs approach VideoQA as a text generation task conditioned on the question $\mathcal{Q}$ and video $\mathcal{V}$ with $T$ frames. Similar to other LMMs, InstructBLIP predicts the answer $\mathcal{A}$ by the following steps:
@@
(1) The Vision Transformer Dosovitskiy et al. (2021) in EVA-CLIP Fang et al. (2023) serves as the frozen image encoder to extract embeddings of each frame individually, and obtains $\textbf{E}=\left\{\textbf{e}_{1},\textbf{e}_{2},\cdots,\textbf{e}_{T}\right\},\textbf{E}\in\mathbb{R}^{T\times N_{I}\times D_{I}},\textbf{e}_{t}\in\mathbb{R}^{N_{I}\times D_{I}}$, where $t$ denotes the $t$-th frame, $N_{I}$ is the patch number of each frame (including the class token), and $D_{I}$ is the embedding dimension. To mitigate computational costs, existing LMMs uniformly sample $K$ frames ($K\ll T$) to represent the video, resulting in the $\hat{\textbf{E}}\in\mathbb{R}^{K\times N_{I}\times D_{I}}$.
@@
(2) A trainable Q-former serves as the connection module to bridge the modality gap between vision and language. It takes frame embeddings $\hat{\textbf{E}}$ and outputs a set of fixed-length frame tokens $\textbf{F}=\left\{\textbf{f}_{1},\textbf{f}_{2},\cdots,\textbf{f}_{K}\right\},\textbf{F}\in\mathbb{R}^{K\times N_{C}\times D_{C}},\textbf{f}_{t}\in\mathbb{R}^{N_{C}\times D_{C}}$, where $N_{C}$ is the token number of each frame ($N_{C}\ll N_{I}$, e.g., $N_{C}=32$ and $N_{I}=257$ in InstructBLIP), and $D_{C}$ is the dimension of the connection module.
@@
(3) Each $\textbf{f}_{t}$ in F are concatenated together to obtain the flattened $\textbf{F}\in\mathbb{R}^{(K\cdot N_{C})\times D_{C}}$, followed with a fully-connected layer to project F into the LLM’s dimension $D_{L}$. At last, the final projected $\textbf{F}\in\mathbb{R}^{(K\cdot N_{C})\times D_{L}}$ is fed into the frozen LLM (e.g., FLAN-T5 Chung et al. (2022) or Vicuna Zheng et al. (2023)) serving as soft prompts, together with the word embeddings of question $\mathcal{Q}$, to generate the answer text $\mathcal{A}$.
@@
The model is trained by optimizing the trainable parameters $\theta$ of the model $P$ with the autoregressive objective:
$\mathcal{L}_{vqa}=-\sum_{t=1}^{L_{a}}logP_{\theta}(\mathcal{A}_{t}|\mathcal{A}_{<t},\mathcal{V},\mathcal{Q})$ (1)
@@
Figure 2: (a) The overall framework of our method. (b) The process of pseudo-label generation.
4 Method
Figure 2 (a) gives an overview of our framework. After extracting the embeddings of $T$ frames, our GCG module will select the most question-critical $K$ frames $\hat{\textbf{E}}$ from E, as the visual inputs for the LMM. To ensure the selected frames are most relevant to answering the question, the GCG module will be additionally optimized by the pseudo-labels of question-critical moments, resulting in the regression objective $\mathcal{L}_{reg}$ from weakly grounded timestamps, and the contrastive objective $\mathcal{L}_{con}$ aligning the paired description-moment pairs while pushing away unpaired ones.
4.1 Inputs Representation
For video representations, along with the frame embeddings $\textbf{E}\in\mathbb{R}^{T\times N_{I}\times D_{I}}$ extracted in step (1) of Section 3, we also extract the corresponding class tokens $\textbf{E}_{[CLS]}\in\mathbb{R}^{T\times D_{I}}$ from E for further pseudo-labels generation and contrastive grounding. For language representations, we tokenize the question $\mathcal{Q}$ into a sequence of words and then feed them into the text encoder of EVA-CLIP, to get word-level embeddings $\textbf{Q}=\left\{\textbf{q}_{t}\right\}_{t=1}^{L_{q}}\in\mathbb{R}^{L_{q}\times D_{I}}$, where $L_{q}$ denotes the sequence length of the question. We get the embeddings of the fused event description (detailed in 4.2) the same way as Q, but only retain the class token $\textbf{d}_{[CLS]}\in\mathbb{R}^{D_{I}}$ to represent it for further pseudo-labels generation.
Figure 3: We use the Gaussian generator to generate multiple Gaussian functions and obtain weight distributions $p\in\mathbb{R}^{T}$ for each video moment. The Gaussian generator will be optimized by the weakly-supervised regression objective $\mathcal{L}_{reg}$ and contrastive objective $\mathcal{L}_{con}$, along with the fully supervised QA objective $\mathcal{L}_{vqa}$, to discover the most question-critical moments as visual inputs for LMMs.
4.2 Pseudo Labels for Temporal Grounding

### v3

@@ Abstract.
Video Question Answering (VideoQA) aims to answer natural language questions based on the information observed in videos. Despite the recent success of Large Multimodal Models (LMMs) in image-language understanding and reasoning, they deal with VideoQA insufficiently, by simply taking uniformly sampled frames as visual inputs, which ignores question-relevant visual clues. Moreover, there are no human annotations for question-critical timestamps in existing VideoQA datasets. In light of this, we propose a novel weakly supervised framework to enforce the LMMs to reason out the answers with question-critical moments as visual inputs. Specifically, we first fuse the question and answer pairs as event descriptions to find multiple keyframes as target moments and pseudo-labels, with the visual-language alignment capability of the CLIP models. With these pseudo-labeled keyframes as additionally weak supervision, we devise a lightweight Gaussian-based Contrastive Grounding (GCG) module. GCG learns multiple Gaussian masks to characterize the temporal structure of the video, and sample question-critical frames as positive moments to be the visual inputs of LMMs. Extensive experiments on several benchmarks verify the effectiveness of our framework, and we achieve substantial improvements compared to previous state-of-the-art methods.
Video Question Answering, Large Multimodal Models
[…]
@@
Figure 1. We argue that the information in uniformly sampled frames is insufficient for LMMs to answer the question correctly. Therefore, we utilize the fused event description to provide additional weak supervision and generate weight distributions for each video moment. We align the positive description-moment pairs while pushing away negative ones. In this way, we can effectively select question-critical moments for LMMs to reason out the answer.
1. Introduction
@@
Video Question Answering (VideoQA) stands at the forefront of developing intelligent systems that can reason about causal and temporal relations and answer natural language questions in videos, which is an essential manifestation of human intelligence. Despite significant advancements have been made by self-supervised pretraining and transformer-style architectures (Radford et al., 2021; Yang et al., 2021; Wang et al., 2023, 2022; Ye et al., 2023b; Yan et al., 2022) in recent years, VideoQA remains a challenging problem that requires models to comprehensively understand and dynamically align the semantics of both the visual world and natural language.
@@
With the progress in vision-language pre-training techniques (Radford et al., 2021; Li et al., 2022a) and Large Language Models (LLMs) (Touvron et al., 2023; Brown et al., 2020), Large Multimodal Models (LMMs) (Dai et al., 2023; Li et al., 2023b; Liu et al., 2023), as the further development of Large Language Models (LLMs) (Touvron et al., 2023; Brown et al., 2020), have showcased impressive capabilities across various image-language tasks. These LMMs share very similar architecture and paradigms. They first extract visual features with an image encoder, and the encoded features will be sent into a connection module to obtain a set of visual tokens that are in the same feature space as the LLM. Then, the visual tokens are concatenated with the input text embeddings together, to be fed into the LLM to decode the target text sequence. However, limited by the long sequence frames in videos and computation costs, current LMMs fall short when applied to VideoQA. They simply concatenate the visual tokens of uniformly sampled, sparse frames (e.g., 4 frames) as the visual inputs for answer prediction. Such a sampling strategy does not consider the specific question at hand, treating all frames equally and introducing redundancy, potentially distracting the model from discovering true answers.
@@
Therefore, it’s necessary to localize the moments crucial for answering the question for LMMs (as the positive moments shown in Figure 1). Notably, different from the task of Temporal Sentence Grounding (TSG) (Gao et al., 2017; Zhang et al., 2023) which aims to localize a video moment described by a declarative sentence, the grounding mechanism in VideoQA features some unique challenges. First, questions in VideoQA are interrogative sentences, and they lack explicit information about the answer content needed to be grounded. For instance, in Figure 1, there is a semantic gap between the interrogative question [How does the boy react after opening the present?] and the declarative description [The boy carries up the dog after opening the present.]. Thus, models are required to both localize the moment [after opening the present] explicitly shown in the question and identify the implicit answer moment [carry up the dog], demanding the causal-temporal reasoning. Second, VideoQA aims to correctly answer the questions of videos, rather than solely grounding specific video moments, and there are no human annotations for the timestamps of question-critical moments in existing VideoQA datasets.
@@
To address these challenges, we introduce a weakly supervised framework by discovering question-critical moments with Gaussian-based Contrastive Grounding (GCG). As labeling the timestamps of question-critical moments is labor-intensive and subjective, we leverage the powerful visual-language alignment capability of the CLIP models (Radford et al., 2021; Fang et al., 2023) to provide timestamps of keyframes. In detail, we fuse the textual question and answer to generate a declarative sentence as the event description, and then compute the similarities between the description and each frame. Frames with the highest scores will be the keyframes of target moments. We observe that LMMs with these pseudo-labeled keyframes as visual inputs showcased significant improvements on a wide range of VideoQA tasks (as shown in Figure 4), which also indicates a great potential to localize question-critical moments for LMMs. To equip LMMs with such ability to automatically find these question-critical moments, motivated by more recent research (Xiao et al., 2023; Zheng et al., 2022a, b) which has highlighted the superiority of end-to-end Gaussian mask learning in video grounding tasks, we use multiple Gaussian masks to characterize the inherent temporal structure of the video. Differently, we explicitly introduce additional objectives as weak supervision to help generate more suitable Gaussians for LMMs. With this new design, our GCG will distinguish the positive video moments (green in Figure 1) from negative video moments (orange and yellow in Figure 1). The positive moments are crucial for answering the question and will be the visual inputs of LMMs for answer prediction. Moreover, to ensure the selected positive moments are closest to the event description, our GCG also includes a contrastive objective (He et al., 2020) that can align the positive description-moment pairs while pushing away negative ones. Notably, different from previous works like SeViLA (Yu et al., 2023) which pre-train an additional LMM as the keyframe localizer with other datasets like QV-Highlights (Lei et al., 2021), our method is lightweight and flexible for end-to-end training with LMMs.
To summarize, we make the following contributions: (1) We propose a weakly supervised grounding framework for VideoQA, by utilizing the alignment capability of the CLIP models to provide pseudo-labeled timestamps of keyframes without human-labor annotated costs. (2) We devise the Gaussian-based Contrastive Grounding (GCG) for weakly-grounded selection of question-critical moments, enhancing the effectiveness and interpretability of LMMs when applied to VideoQA, by revealing which visual scenes result in the predicted answers. (3) We conduct extensive experiments to verify the effectiveness of our proposed method, and achieve substantial improvements on six challenging VideoQA benchmarks including NExT-QA, Causal-VidQA, Intent-QA, ActivityNet-QA, MSVD-QA, and MSRVTT-QA.
2. Related Works
2.1. Large Multimodal Models (LMMs)
@@
LMMs (Dai et al., 2023; Li et al., 2023b; Ye et al., 2023a; Liu et al., 2023; Alayrac et al., 2022) in their current form primarily function as image-to-text generative models, taking images as input and generating text sequences. These models have demonstrated strong capabilities in image-language understanding and reasoning by adapting frozen language models to frozen image encoders with trainable connection modules, following large-scale image-text pretraining. The connection module can either be a transformer-based architecture like Q-former in InstructBLIP and BLIP-2 (Dai et al., 2023; Li et al., 2023b), Perceiver Resampler in Flamingo (Alayrac et al., 2022), or a simple linear layer in LLaVA (Liu et al., 2023). Most current LMMs are essentially image-based models, and they simply concatenate the visual tokens extracted from uniformly sampled, sparse frames as visual inputs for video-language tasks. This results in a lack of temporal modeling ability and emphasizes the necessity of selecting specific video moments, particularly for addressing the demands of reasoning-based VideoQA tasks. In this paper, our goal is to enhance the causal-temporal reasoning abilities of LMMs without additional pretraining on video-text corpora, by discovering the question-critical moments with our weakly-supervised Gaussian-based Contrastive Grounding.
2.2. Temporal Grounding in VideoQA
@@
Early VideoQA benchmarks (Yu et al., 2019; Xu et al., 2017; Jang et al., 2017) focus on descriptive questions (e.g., [what’s the man doing]) within short video clips, rarely going beyond a recognition of the objects and actions. Instead, more recent VideoQA benchmarks (Li et al., 2022b; Xiao et al., 2021; Li et al., 2023d) like NExT-QA (Xiao et al., 2021) emphasize counterfactual, temporal, and causal reasoning involving multiple entities and relations, demanding the ability to uncover the causes behind specific events within longer videos, necessitating localizing a text query to specific moments.
@@
In light of this, ATP (Buch et al., 2022) utilizes the tool of atemporal probe to select a single frame without temporal information for downstream tasks. MIST (Gao et al., 2023) and TranSTR (Li et al., 2023e) fuse frames with the mechanism of adaptive temporal rationalization and iterative spatial-temporal attention, respectively. NExT-GQA (Xiao et al., 2023) constructs grounding labels in the test set of NExT-QA and uses a single Gaussian mask to fuse frames along the temporal dimension. SeViLA (Yu et al., 2023), similar to us, utilizes the LMM (BLIP-2) for VideoQA. However, SeViLA uses two LMMs to generate pseudo-labels and answer questions respectively, with extra pre-training on TSG datasets (Lei et al., 2021) and a multi-stage training scheme. Different from previous works, we utilize the CLIP (Radford et al., 2021; Fang et al., 2023) models to automatically provide weak supervision for grounding, and our lightweight GCG module learns multiple Gaussian masks to generate both positive and negative moments in an end-to-end manner, with an additionally contrastive objective to distinguish positive ones from negative ones for frame selection.
3. Preliminary: LMMs for VideoQA
@@
We take InstructBLIP (Dai et al., 2023) as an example to illustrate how LMMs deal with VideoQA. Similar to other LMMs, InstructBLIP approaches VideoQA as a text generation task conditioned on the question $\mathcal{Q}$ and video $\mathcal{V}$ with $T$ frames, and predicts the answer $\mathcal{A}$ by the following three steps:
@@
(1) The Vision Transformer (Dosovitskiy et al., 2021) in EVA-CLIP (Fang et al., 2023) serves as the frozen image encoder to extract embeddings of each frame individually, and obtains $\textbf{E}=\left\{\textbf{e}_{1},\textbf{e}_{2},\cdots,\textbf{e}_{T}\right\},\textbf{E}\in\mathbb{R}^{T\times N_{I}\times D_{I}},\textbf{e}_{t}\in\mathbb{R}^{N_{I}\times D_{I}}$, where $t$ denotes the $t$-th frame, $N_{I}$ is the patch number of each frame (including the class token), and $D_{I}$ is the embedding dimension. To mitigate computational costs, existing LMMs uniformly sample $K$ frames ($K\ll T$) to represent the video, resulting in the sampled $\hat{\textbf{E}}\in\mathbb{R}^{K\times N_{I}\times D_{I}}$.
@@
(2) A trainable Q-former serves as the connection module to bridge the modality gap between vision and language. It takes frame embeddings $\hat{\textbf{E}}$ as inputs and outputs a set of fixed-length frame tokens $\textbf{F}=\left\{\textbf{f}_{1},\textbf{f}_{2},\cdots,\textbf{f}_{K}\right\},\textbf{F}\in\mathbb{R}^{K\times N_{C}\times D_{C}},\textbf{f}_{t}\in\mathbb{R}^{N_{C}\times D_{C}}$, where $N_{C}$ is the token number of each frame ($N_{C}\ll N_{I}$, e.g., $N_{C}=32$ and $N_{I}=257$ in InstructBLIP), and $D_{C}$ is the dimension of the connection module.
@@
(3) Each $\textbf{f}_{t}$ in F are concatenated together to obtain the flattened $\textbf{F}\in\mathbb{R}^{(K\cdot N_{C})\times D_{C}}$, followed with a fully-connected layer to project F into the LLM’s dimension $D_{L}$. At last, the final projected $\textbf{F}\in\mathbb{R}^{(K\cdot N_{C})\times D_{L}}$ is fed into the frozen LLM (e.g., FLAN-T5 (Chung et al., 2022) or Vicuna (Zheng et al., 2023)) serving as soft prompts, together with the word embeddings of question $\mathcal{Q}$, to generate the answer text $\mathcal{A}$.
@@
The model is trained by optimizing the trainable parameters $\theta$ of the model $P$ with the autoregressive language modeling objective:
(1) $\mathcal{L}_{vqa}=-\sum_{t=1}^{L_{a}}logP_{\theta}(\mathcal{A}_{t}|\mathcal{A}_{<t},\mathcal{V},\mathcal{Q})$
@@
4. Method
Figure 2 (a) gives an overview of our framework. After extracting the frame embeddings E as in Section 3, step (1), our GCG will select the most question-critical $K$ frames $\hat{\textbf{E}}$ from E, as the visual inputs for the LMM. To ensure the selected frames are most relevant to answering the question, the GCG module will be additionally optimized by the pseudo-labels of question-critical moments, resulting in the regression objective $\mathcal{L}_{reg}$ from weakly grounded timestamps, and the contrastive objective $\mathcal{L}_{con}$ aligning the paired description-moment pairs while pushing away unpaired ones.
Figure 2. (a) The overall framework of our method. (b) The process of pseudo-label generation.
4.1. Inputs Representation
For video representations, along with the frame embeddings E, we also extract the corresponding class tokens $\textbf{E}_{[CLS]}\in\mathbb{R}^{T\times D_{I}}$ from E for further pseudo-labels generation and contrastive grounding. For language representations, we tokenize the question $\mathcal{Q}$ into a sequence of words and then feed them into the text encoder of EVA-CLIP, to get word-level embeddings $\textbf{Q}=\left\{\textbf{q}_{t}\right\}_{t=1}^{L_{q}}\in\mathbb{R}^{L_{q}\times D_{I}}$, where $L_{q}$ denotes the sequence length of the question. We get the embeddings of the fused event description (detailed in 4.2) the same way as Q, but only retain the class token $\textbf{d}_{[CLS]}\in\mathbb{R}^{D_{I}}$ to represent it for further pseudo-labels generation.
Figure 3. We use the Gaussian generator to generate multiple Gaussian masks and obtain weight distributions $p\in\mathbb{R}^{T}$ for each video moment. The Gaussian generator will be optimized by the regression objective $\mathcal{L}_{reg}$ and contrastive objective $\mathcal{L}_{con}$, along with the fully supervised QA objective $\mathcal{L}_{vqa}$, to discover the most question-critical moments as visual inputs for LMMs.
4.2. Pseudo Labels for Temporal Grounding